OpenStudy (anonymous):

The LCM of 9 and 12 is _______. I WILL GIVE MEDAL N BE A FAN!!!<<<<<<

OpenStudy (gorv):



OpenStudy (gorv):

to find lcm we divide our number with prime number can u tell me smallest + prime number

OpenStudy (gorv):

i will give a hint 2,3,5,7

OpenStudy (gorv):

now??

OpenStudy (igreen):

LCM is the smallest number that can get into 9 and 12(not including one) So what do you think it is?

OpenStudy (anonymous):

OK HOLD ON IM DOING IT ON PAPER

OpenStudy (igreen):

Okay.

OpenStudy (gorv):

okay give it a try

OpenStudy (anonymous):

1

OpenStudy (anonymous):

12 = 1 * 2 * 2 * 3; 9 = 1 * 3 * 3. I have bolded the factors in common. Multiply the factors in common by the other factors: 1 * 3 * 2 * 2 * 3 = 36, so the LCM of 9 & 12 is 36

OpenStudy (igreen):

Lol that's wrong @calebwill123

OpenStudy (igreen):

36 can't get into 9 and 12 because it's way too big.

OpenStudy (gorv):

after 1 nest prime no is

OpenStudy (anonymous):

oh thx @iGreen

OpenStudy (gorv):

okk prime number are the number which are divisible by 1 or itself so next is 2

OpenStudy (anonymous):

OK

OpenStudy (gorv):

so we start with 2 generally

OpenStudy (gorv):

so in 9 and 12 is there any no which can be divided by 2 ??

OpenStudy (anonymous):

YES

OpenStudy (igreen):

You don't want any decimals.